# Photoluminescence from metal nanostructures - dependence on size

### v1

@@ 3.2.3 Comparison to experiments
As done above for single particles, we now compare the predictions of our simple analytic formula to recent experimental measurements. Specifically, we look at the data of Bowman et al. 28 Fig. 3(b) and Fig. S20(b), which presents the ratios of PL measurements from single crystal Au layers of various thicknesses to that of the thickest layers. Only a few layer thicknesses were studied, but data is available for several emission frequencies (rather than for single frequencies, as for spheres).
The normalized PL spectra presented in Fig. 3(c) and Fig. S20(b) from Bowman et al. 28 are reproduced in Fig. 8(a) and (c) and Fig. 8(b) and (d), respectively, along with our own theory of PL of films (Eq. (9)). Our calculations closely match the experimental and theoretical results of Bowman et al. 28, for all thicknesses studied (see results for the thicker films nm and nm in Fig. 17). There are, however, small discrepancies observed for thinner films occurring at different parts of the spectra. In order to check whether they represent a quantum effect, as claimed in Ref. [ 28], the PL is also calculated by modifying the Drude damping constant to , where is the Fermi velocity and is the proportionality factor that can range from to 69. We show that the choice yields an improved match for the excitation at nm, but for the PL calculated for excitation at nm it actually causes a deviation from the experimental data.
In light of this overall good agreement, it is notable that our simple theory deviates only slightly from the sophisticated quantum mechanical calculations which were highlighted in the original work 28; indeed, the theory in that study involved the summation of the individual dipole emissions, a treatment of the electron states as a discrete set in momentum space, and a detailed DFT-based calculation of the permittivity. Those we replaced by the local Kirchhoff Law, empiric permittivity data, and our simple analytic formula for PL (Eq. (9)) based on a continuous set of energy states, along with standard macroscopic electromagnetic and thermal calculations. In that sense, beside the inconclusive role of the quantum size effect, the only quantum aspect of our theory is the use of the quantum version of the Boltzmann equation (see Ref. [ 39]), while the electron states themselves (and thus, the permittivity) nor the photon properties do not involve quantization. In fact, our theoretical predictions match the more sophisticated theory of Ref. [ 28] even where the latter does not match the experimental data well.
@@ 4 Discussion
In this work, we provide a unifying explanation for the dependence of the PL from metal nanostructures on their size. Specifically, using the experimentally-established expression for the non-equilibrium electron distribution in the conduction band 39, 40, 42, 43, 44, together with recent extension of the local Kirchhoff Law for light emission from metals characterized this non-equilibrium electron distribution 36, we provide a simple analytical form for the (intensity, temperature and) size-dependence of the PL. In particular, we show that the size-dependence of the Stokes emission (SE) is determined primarily by the size-dependence of the absorption (= emission) cross-section density (aka emissivity density) but that the aSE is much more sensitive to the structure size at high illumination intensities due to the exponential dependence of the Bose function on the electron temperature .
@@
In particular, our predictions imply that, generically, the PL from spheres and layers grows for small sizes as the volume / thickness, respectively, and then decreases due to the decrease of the average local field (or equivalently, of the emissivity). The size for which maximal PL is attained is found to be larger for spheres than for layers (e.g., nm vs. nm, respectively, for illumination near resonance; the maximal size grows for off-resonant illumination). This explains the seemingly contradicting trends observed (e.g., Ref. [ 23, 24, 16]) - an increase of the PL from spheres of growing size, and a decrease of the PL from layers of growing thickness. Specifically, our analysis shows that the different size-dependence of the PL reported in experiments originates from the fact that measurements from thin films 16, 28 were so far performed only for layers thicker than the size of maximal PL ( nm), most likely, because films of such thickness tends to be discontinuous (below the percolation limit). A similar behaviour was observed for rods.
@@
Our predictions were shown to match well a range of experimental results of SE, including for cases that formally lie beyond the limits of validity of the analysis, including non-spherical particles and their arrays. The quantitative discrepancies observed in some cases may originate in differences in the underlying permittivity data, in subtleties associated with the significant field non-uniformity in this case, or in missing elements needed for a complete description of the PL, e.g., the inclusion of emission originating from inelastic electron scattering, which was already shown to dominate the emission near the pump frequency 32. Even better understanding of these discrepancies may be obtained by comparing to aSE measurements, which were so far done in the context of thermometry 5, 6, 8, 7, but not done systematically as a function of the size.
@@
Our analytic approach provides an easily accessible alternative to the highly complicated rigorous (discrete -space) calculations of the PL offered in some recent studies 29, 28, 32. In particular, the use of the extension of the local Kirchhoff Law to non-equilibrium electron distributions 36 allows circumventing many of the detailed electromagnetic emission calculations done in Ref. [ 28] and validates approximations made, e.g., the neglect of the energy-dependence of the matrix elements and electron density of states; this finding is in line with good agreement obtained when matching our theoretical approach to measurement of currents in molecular and current junctions 42, 41, 43, 44. Moreover, while the analytic expression we provide accounts for both intraband and interband absorption events (via the empiric values of the permittivity), unlike the rigorous approaches mentioned, our approach accounts only for emission events occurring within the conduction band. Nevertheless, our predictions are found to be in good agreements with experimental data from Au nanostructures in regimes for which interband emission was shown to be significant. This implies that interband emission events have primarily a quantitative effect on the PL, but not a qualitative one.
@@
Nevertheless, while both rigorous -space approach and our approximate analytic approaches are capable of reproducing most aspects of the experimental observations, the current study highlights the ability of the more rigorous approaches to capture subtle aspects such as the determination of the conditions under which the emission is due to radiative recombination or inelastic scattering (see, e.g., discussions in Ref. [ 74, 75, 76, 77, 78, 4, 32, 28]), and identify the mechanism underlying the high sensitivity of the PL to the number of atomic levels observed in Ref. [ 16].
While in the current work we focused on the PL from simple structures and noble metals, our approach can be applied also to more complicated structures for which the local electric field is even more non-uniform such as long rods 79, various particle dimers 80, 79, 32, 81, trimers 82 etc., or to other plasmonic materials, including low electron density Drude materials such as transparent conducting oxides 83, 84.
@@
We also emphasize that the analysis in the current manuscript is limited to CW illumination, but could be extended also to pulsed illumination 9, 20, 10, 16, 85. In this scenario, which was studied more extensively in the past, the size-dependence of the temperature is different, and there is a dynamic transition between the non-thermal and thermal parts (see Ref. [ 34]), which entails further complexity. This class of experiments will be analyzed separately in future work.
Ultimately, our work paves the way to simple optimization of the PL for practical purposes.
@@
Acknowledgements. I. K. and Y.S. were partially funded by Israel Science Foundation (ISF) grant no. 340/2020, a Lower-Saxony - Israel collaboration grant no. 76251-99-7/20 (ZN 3637) as well as a grant from the Bronitzki family.
Appendix A Appendix: Additional figures

### v2

@@ 3.2.3 Comparison to experiments
As done above for single particles, we now compare the predictions of our simple analytic PL formula to recent experimental measurements. Specifically, we look at Figs. 3(a)-(c) and Figs. S20(a)-(b) of Bowman et al. 29, which presents PL measurements from single crystal Au films of various thicknesses. Only a few film thicknesses were studied, but data is available for several emission frequencies (rather than for single frequencies, as for particles). The PL recorded for the individual films is shown in Fig. 8, alongside our calculations. To conform with the data of the original paper, our results for nm (Fig. 8(a) and (c)) are given in units of nm-1 (as in Figs. 3(a)-(c) of Ref. [ 29], multiplied by a constant factor of 8), while those for nm (Fig. 8(b) and (d)) are presented in arbitrary units (as in Fig. S20(a) of Ref. [ 29], and are normalized using a method analogous to that employed for the experimental results of spheres shown in Fig. 4(a). For nm excitation, the calculated results tend to fall below the experimental data at shorter wavelengths. This discrepancy likely arises from the omission of interband contributions in our emission calculations.
In order to remove the impact of instrumental noise and a substrate-related Raman peak, we follow Ref. [ 29] and show the PL ratio spectra for the same films (i.e., the PL data / difference normalized by the data from a thick film) which are presented in Fig. 3(c) and Fig. S20(b) from Bowman et al. 29 in Fig. 9(a) and (c) and Fig. 9(b) and (d), respectively, along with our own theoretical prediction for films (Eq. (9)). For the nm excitation, despite the differences observed at short wavelengths in Fig. 8(a) and (c), our calculations now closely match the theoretical results of Bowman et al. 29 for all thicknesses studied (see results for the thicker films (see Fig. LABEL:Fig_exp_Bowman_thicker). A similar match can be observed for nm excitation as well (see Fig. 9(b) and (d)). In fact, our theoretical predictions match the more sophisticated theory of Ref. [ 29] even in the parts of the spectra where the latter does not match the experimental data well, leaving the discrepancies unexplained by either theoretical approach. This is remarkable, since the theory in Ref. [ 29] involved the summation of the individual dipole emissions, a treatment of the electron states as a discrete set in momentum space, and a detailed DFT-based calculation of the permittivity. Those we replaced by the local Kirchhoff Law, empiric permittivity data, and our simple analytic formula for the PL (Eq. (9)) based on a continuous set of energy states, along with standard macroscopic electromagnetic and thermal calculations. Thus, overall, the quantitative match for the normalized data implies that the contribution of interband transitions to the PL exhibits a size dependence very similar to the intraband transitions -based prediction.
@@ 4 Discussion
In this work, we provide a unifying explanation for the dependence of the PL from metal nanostructures on their size. Specifically, using the experimentally-established expression for the non-equilibrium electron distribution in the conduction band 40, 41, 43, 44, 45, together with recent extension of the local Kirchhoff Law for light emission from metals characterized this non-equilibrium electron distribution 37, we provide a simple analytic form for the (intensity, temperature and) size-dependence of the PL. In particular, we show that the size-dependence of the Stokes emission (SE) is determined primarily by the size-dependence of the absorption (= emission) cross-section density (aka emissivity density) but that the aSE is much more sensitive to the structure size at high illumination intensities due to the exponential dependence of the Bose function on the electron temperature .
@@
This behaviour is illustrated schematically in Fig. 10. As seen, generically, the PL from particles and films grows for small sizes as the volume / thickness, respectively, and then decreases due to the decrease of the average local field (or equivalently, of the emissivity). The size for which maximal PL is attained is found to be larger for particles than for films ( nm vs. nm, respectively, for illumination near resonance; the maximal size is higher for off-resonant illumination). This explains the seemingly contradicting trends observed in experiments (e.g., Ref. [ 24, 25, 17]) - they originate from the fact that measurements from thin films 17, 29 were so far performed only for films thicker than the size of maximal PL ( nm), most likely, because thinner films tend to be discontinuous (i.e., below the percolation limit). A similar behaviour was observed for rods. Once advanced fabrication of sub 10nm thick films will be employed for (CW) PL studies (as, e.g., in Ref. [ 17]), the quantum effects predicted in Ref. [ 29] could be identified. Notably, the trends described in Fig. 10 apply both for spectrally-resolved and spectrally-integrated PL; moreover, the trends will be qualitatively similar also for different wavelengths, for higher temperatures/intensities and for the aSE, in which cases, the sensitivity to the size will be greater.
@@
Our predictions were shown to match well a range of experimental results of SE, including for cases that formally lie beyond the limits of validity of the analysis, including non-spherical particles and their arrays. The quantitative discrepancies observed in some cases may originate in differences in the underlying permittivity data, in subtleties associated with the significant field non-uniformity in this case, or in missing elements needed for a complete description of the PL, e.g., the inclusion of emission originating from inelastic electron scattering555According to Ref. [ 33], this contribution dominates the emission near the pump frequency. , interband transitions and the subtle QM effects highlighted in Ref. [ 29]. Even better understanding of these discrepancies may be obtained by comparing to aSE measurements, which were so far done in the context of thermometry 6, 7, 9, 8, but not done systematically as a function of the structure size.
@@
Our analytic approach provides a very simple alternative to the highly complicated rigorous (discrete -space) calculations of the PL offered in some recent studies 30, 29, 33. In particular, the use of the extension of the local Kirchhoff Law to non-equilibrium electron distributions 37 allows circumventing many of the detailed electromagnetic emission calculations done in Ref. [ 29] and validates the approximations we made, e.g., the neglect of the energy-dependence of the matrix elements and electron density of states; this finding is in line with the good agreement obtained when matching our theoretical approach to measurement of currents in molecular and current junctions 43, 42, 44, 45. Moreover, while the analytic expression we provide accounts for only intraband transitions, it is found to yield good qualitative and sometimes even quantitative agreement with experimental data from Au nanostructures in regimes for which interband emission was shown to be significant666In this context, the modelling of the contribution of interband transitions to the PL in Ref. [ 18] should be questioned, especially since it is based on an unjustified fold reduction of the electron-phonon coupling coefficient. .
@@
Nevertheless, while both rigorous -space approach and our approximate analytic approaches are capable of reproducing most aspects of the experimental observations, the current study highlights the ability of the more rigorous approaches to capture subtle aspects such as the determination of the conditions under which the emission is due to radiative recombination or inelastic scattering (see, e.g., discussions in Ref. [ 76, 77, 78, 79, 80, 5, 33, 29]), emission from pre-scattered electrons 29, identifying subtle quantum mechanical effects that go beyond the standard quantum size effect (e.g., Ref. [ 81, 82])777In comparison, the only quantum aspect of our theory is the use of the quantum version of the Boltzmann equation (see Ref. [ 40]), while the electron states themselves (and thus, the permittivity) and the photon properties do not involve quantization. such as the thickness dependence of the CW PL from atomic flat metal films (see Fig. 4a of Ref. [ 29]) and in the future, maybe even the high sensitivity to the number of atomic levels observed in Ref. [ 17] for the PL due to pulsed illumination.
@@
While in the current work we focused on the PL from simple structures and noble metals, our approach can be applied also to more complicated structures for which the local electric field is even more non-uniform such as long rods 83, various particle dimers 84, 83, 33, 85, trimers 86 etc., or to other plasmonic materials, including low electron density Drude materials such as transparent conducting oxides 87, 88. We also emphasize that the analysis in the current manuscript is limited to CW illumination, but could be extended also to pulsed illumination 10, 21, 11, 17, 89. In this scenario, which was studied more extensively in the past, the size-dependence of the temperature is different, and there is a dynamic transition between the non-thermal and thermal parts (see Ref. [ 35]), which entails further complexity. This class of experiments will be analyzed separately in future work. Ultimately, our work paves the way to simple optimization of the PL for practical purposes.
@@
Acknowledgements. I. K. and Y.S. were partially funded by Israel Science Foundation (ISF) grant no. 340/2020, a Lower-Saxony - Israel collaboration grant no. 76251-99-7/20 (ZN 3637) as well as a grant from the Bronitzki family. I.W.U. was funded by the Guangdong Natural Science Foundation (Grants No. 2024A1515011457).
The Supporting Information is available free of charge.
Absorption cross-section of spheres, absorption characteristics across spectrum, thermal and non-thermal contribution to PL of sphere, PL of sphere excited off-resonance, PL calculated for higher order mode responses, the thermo-optic effect on absorption, PL calculation for films.